OpenStudy (anonymous):

1. which inequality matches the graph? 2n + 5 < 7 (this < sign has a line under it) 2n - 5 < 7 2n + 5 < 7 2n + 5 > 7 (this > sign has a line under it) 2. which inequality matches the graph? 6w + 6> 18 6w + 6> 18(this > sign has a line under it) 6w - 6> 18 -6w + 6> 14

OpenStudy (anonymous):

this is for Q1

OpenStudy (anonymous):

this is for Q2

OpenStudy (anonymous):

@GIitter

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@GIitter

OpenStudy (anonymous):

@soprano.h.d0816

imqwerty (imqwerty):

alright for 1st one we got this- and the options are these- \(2n + 5 \le 7\) \(2n-5<7\) \(2n+5<7\) \(2n+5 \ge 7\) 1st thing to notice is-> \(n\) is nothing but the values the graph can obtain. 1st we analyze the graph well in the graph all values starting from 1 and going to \(-\infty\) are included so we can say that -> all values are less than or equal to 1. so basically-> \(n \le 1\) now we look at our options and try to simplify them so that we can obtain this form-> \(\color{blueviolet}{n \le 1}\) alright so we start from option 1 and simplify it \(2n +5 \le 7\) \(2n+5-5 \le 7-5\) \(2n \le 2\) \(\large \frac{2n}{2} \le \frac{2}{2}\) \(n \le 1\) :) this is what we wanted so in question 1 option 1 is correct the second problem is similar try to do that (:

OpenStudy (anonymous):

ok so the first question is A

imqwerty (imqwerty):

yeah =]

OpenStudy (anonymous):

cool what about 2

imqwerty (imqwerty):

you wanna try it urself? (B

OpenStudy (anonymous):

ok i'll try:/

imqwerty (imqwerty):

okay lol if u get stuck you can always ask for help :)

OpenStudy (anonymous):

c?

imqwerty (imqwerty):

noo you are close tho

OpenStudy (anonymous):

oh

imqwerty (imqwerty):

:D

OpenStudy (anonymous):

A

imqwerty (imqwerty):

\(\Huge\ddot\smile\) correct!!

OpenStudy (anonymous):

i have one more question can you help me real quick with it

imqwerty (imqwerty):

alright :)

OpenStudy (anonymous):

An airline requires carry-on luggage to weigh at most 40 pounds. Your suitcase currently weighs 10 pounds. How many pounds p are available for you to fill your suitcase with other items? p - 10 ≤ 40; p ≤ 50 40 - p < 10; p < 30*** [think it is this one] p + 10 ≥ 40; p ≥ 30 10 + p ≤ 40; p ≤ 30

OpenStudy (anonymous):

close this question open another

OpenStudy (anonymous):

no i am not

OpenStudy (anonymous):

it is in the code of conduct you can be reported for that

imqwerty (imqwerty):

:) its alright since you are new and don't know next time please make another post for another question like 1 question per post or 2 at most :)

OpenStudy (anonymous):

omg seresly

OpenStudy (anonymous):

not u imqwerty i was talking to gigiheart

OpenStudy (anonymous):

sorry I did not know you were new

imqwerty (imqwerty):

well if you ask many question on a single post it might happen that the post starts lagging due to so many comments there. So they recommend making new posts for new questions anyways :) back to our problem alright so we can have 40 pounds at most we already have 10 pounds so if we wanna add \(p\) more pounds of luggage to our bag then- so \(p\) can be 30 pounds at max anyways the total weight of our luggage must be less than or equal to 40 so \(p+10 \le 40\) ; \(p \le 30\)

OpenStudy (anonymous):

thank you so much

imqwerty (imqwerty):

np :)

OpenStudy (anonymous):

wait they dont have that answer choice

OpenStudy (anonymous):

p + 10 ≥ 40; p ≥ 30 there is this way

imqwerty (imqwerty):

see the last one :)

OpenStudy (anonymous):

you but the arrows the other way

imqwerty (imqwerty):

i mean last option

OpenStudy (anonymous):

oh ok

imqwerty (imqwerty):

(:

OpenStudy (anonymous):

thx i got only one wrong but its not the once you helped me with

imqwerty (imqwerty):

np

OpenStudy (anonymous):

i gave you a medal because I saw you tried goodjob and goodluck for next question :D♥♥

OpenStudy (anonymous):

thx